OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6.

OpenStudy (anonymous):

y+6=-1/3(x-6) y+6=-x/3 +2 y+ -x/3 -4

OpenStudy (anonymous):

y= -x/3 -4